# Learning Primitive Embodied World Models: Towards Scalable Robotic Learning

### v2

@@ Abstract
While video-generation-based embodied world models have gained increasing attention, their reliance on large-scale embodied interaction data remains a key bottleneck. The scarcity, difficulty of collection, and high dimensionality of embodied data fundamentally limit the alignment granularity between language and actions and exacerbate the challenge of long-horizon video generation—hindering generative models from achieving a "GPT moment" in the embodied domain. There is a naive observation: the diversity of embodied data far exceeds the relatively small space of possible primitive motions. Based on this insight, we propose a novel paradigm for world modeling–Primitive Embodied World Models (PEWM). By restricting video generation to fixed short horizons, our approach 1) enables fine-grained alignment between linguistic concepts and visual representations of robotic actions, 2) reduces learning complexity, 3) improves data efficiency in embodied data collection, and 4) decreases inference latency. By equipping with a modular Vision-Language Model (VLM) planner and a Start-Goal heatmap Guidance mechanism (SGG), PEWM further enables flexible closed-loop control and supports compositional generalization of primitive-level policies over extended, complex tasks. Our framework leverages the spatiotemporal vision priors in video models and the semantic awareness of VLMs to bridge the gap between fine-grained physical interaction and high-level reasoning, paving the way toward scalable, interpretable, and general-purpose embodied intelligence.
@@ 1 Introduction
Figure 1: While densely distributed data can enable broad generalization, embodied data often suffer from sparsity (Du & Kaelbling, 2024; Xue et al., 2025b). The rightmost schematic highlights how organizing embodied data at the primitive level–along orthogonal dimensions such as action and object–supports compositional generalization even under limited data availability.
Embodied agents capable of planning and decision-making in complex environments rely heavily on internal representations of the world, commonly referred to as world models. Recent advances in pretrained video generation models–powered by self-supervised learning on vast internet-scale datasets–have demonstrated remarkable capabilities in synthesizing fine-grained, temporally coherent visual sequences from natural language descriptions (Du et al., 2023).
These successes have inspired a growing body of work to leverage such models as world models for embodied agents, casting policy learning as a video generation problem (Ko et al., 2023; Du et al., 2023; Yang et al., 2023; Brooks et al., 2024; Black et al., 2024a; Xiang et al., 2024; Bruce et al., 2024; Hu et al., 2024; Zhou et al., 2024; Dalal et al., 2025; Liang et al., 2024a; Clark et al., 2025; Hu et al., 2024; Li et al., 2025; Team et al., 2025a; Pertsch et al., 2025; Zhen et al., 2025b; c; a; Zhu et al., 2025; Hafner et al., 2025; Sun et al., 2024b; a; Ye et al., 2024a): given a language-specified goal, the model generates a future video roll-out, from which actions are subsequently extracted. This paradigm offers compelling advantages in interpretability, generalization, and unified representation across diverse environments.
@@ 2.1 The Motivation and Definition of Primitive Embodied Data
Collecting real-world embodied data is labor-intensive. Existing datasets, despite size, suffer from limited generalization beyond minor scene variations (Xue et al., 2025b; Brohan et al., 2022; Collaboration et al., 2023; Khazatsky et al., 2024b; Walke et al., 2023; Dalal et al., 2025).
To improve generalization in embodied agents, we advocate organizing and collecting data at the level of semantically grounded primitives:
Definition 2.1 (Primitive as Irreducible Generator).
A primitive is a non-constant, finite-duration trajectory $p\in C^{0}([0,T],X)$, $T>0$, that is indecomposable under concatenation: there do not exist $\xi_{1},\xi_{2}\in C^{0}([0,\cdot],X)$ with $0<\mathrm{dom}(\xi_{1}),\mathrm{dom}(\xi_{2})<T$ such that $p=\xi_{1}\cdot\xi_{2}$, and such that $p$ is minimal in the generating set of the behavioral semigroup $\mathcal{S}\subset\mathcal{T}(X)$.
This is grounded in a simple yet powerful observation: the diversity of embodied behaviors vastly exceeds the space of feasible primitive motions. Formally, we posit:
Assumption 2.2 (Finite Primitive Basis).
The set of motion primitives $\mathcal{P}$ is finite, and the space of embodied behaviors $\mathcal{D}_{\text{emb}}$ is approximated arbitrarily well by compositions of elements from $\mathcal{P}$.
A full foundation to derive Assumption 2.2 from minimal empirical axioms is presented in Appendix A. This rationale is also illustrated in Figure 1, which presents a heuristic visualization of the impact of primitive structuring on generalization. Figure 2 illustrates a quick example.
Our design advances world model learning in two orthogonal ways: 1) enabling focused, semantically aligned primitive learning and 2) supporting compositional generalization. Short-horizon primitives reduce temporal complexity, helping the model capture key interaction dynamics with greater precision. Their atomic structure allows complex behaviors to be composed from simple elements, enhancing generalizability. Additionally, the semantic granularity simplifies labeling via vision-language models, while short sequences lower computational and optimization burdens in diffusion training.
2.2 Data Collection: Enhancement in Efficiency, Density, and Quality
@@
We construct a primitive-centric dataset $\mathcal{D}_{\text{prim}}=\{E_{i}\}$, where each episode $E_{i}$ is decomposed into primitives $\{\mathcal{P}^{k}\}_{k=1}^{N_{i}}$ with $\mathcal{P}^{k}=(\mathbf{x}_{t^{k}}^{\text{img}},\text{Instr}^{k},H^{k}_{s\to g},\mathbf{x}_{t^{k}+1:t^{k+1}-1}^{\text{img}})$. Here, $\mathbf{x}_{t^{k}}^{\text{img}}$ is the initial image, $\text{Instr}^{k}$ the instruction, $H^{k}_{s\to g}$ the start-goal heatmap, and future frames $\mathbf{x}_{t^{k}+1:t^{k+1}-1}^{\text{img}}$ are to be predicted.
@@
2.3 Full-Arm Shotting Facilitating Learning-Based Embodied World Simulator
Another notable design choice in $\mathcal{D}_{\text{prim}}$ is that the full robot arm–including the base and joint structures–is meant to be visible in the field of view, in contrast to cropped, end-effector-only frames common in prior datasets. This enables the diffusion model to internalize soft physical constraints such as reachability, joint limits, and spatial feasibility during prediction.
@@ 2.3 Full-Arm Shotting Facilitating Learning-Based Embodied World Simulator
2.4 The High Potential of Leveraging Action-Free Video Data for Scalable, Web-Scale Embodied Learning
It is worth noting that the LIBERO (Liu et al., 2024a) data we use not only includes the original dataset’s multi-view replays but also videos generated through OpenVLA rollouts as training data. This directly demonstrates our method’s ability to leverage any embodied video–whether from rollouts of other policies or from web-scale data–to improve model performance, highlighting one of the key advantages of using video as a universal representation.
@@ 3.1 Dual-Level Compositional Generalization
Implicit Compositionality of Diffusion Video Generation Diffusion models are capable of generating photo-realistic images that combine elements which likely do not appear together in the training set, demonstrating the ability to compositionally generalize (Liu et al., 2022a; Liang et al., 2024b).
@@
We ground this compositional generalization in an Energy-Based Model (EBM) perspective. EBMs naturally support compositionality through additive energy decomposition:
$E(\mathbf{x})=\sum_{i}E_{i}(\mathbf{x};\phi_{i}),$
where each term $E_{i}$ corresponds to a semantic factor (e.g., object, action, or spatial relation). Novel combinations arise by recombining known factors via energy minimization.
@@
Given the diverse and realistic co-occurrence of embodied concepts–such as shape, motion, and spatial relations–in our training data, the model learns rich semantic priors within each 32-frame primitive, enabling strong compositional generalization. Spatiotemporal coherence is inherited from large-scale video pretraining, a key advantage over static image-based methods. We treat each primitive as a distribution and use a heatmap as conditional input, formulating generation as:
$\mathbf{x}_{1:T}^{\text{img}}\sim P(\mathbf{x}_{1:T}^{\text{img}}\mid H_{s\to g},\text{img}_{0}).$
We adopt an I2V model as the world model $\mathcal{W}:(\mathbf{x}^{\text{img}}_{0},H_{s\to g})\mapsto\mathbf{x}^{\text{img}}_{1:T}$. We finetune on DynamiCrafter, which is trained on data without robotic arms. Our method performs well on robotic tasks, demonstrating effective zero-shot transfer. See Appendix F.1 for details.
Explicit Compositionality by Sequentially Combining Primitives By sequentially composing primitives, we treat the Primitive-Enabled World Model (PEWM) as a plug-and-play module for constructing complex behaviors. Each primitive $\pi_{i}\in\mathcal{P}$ represents a reusable, low-level dynamic policy that maps a current image $\mathbf{x}_{0}$ and a goal heatmap $h_{i}$ to a future video segment:
$\mathbf{x}_{1:T_{i}}\sim\mathcal{W}(\mathbf{x}_{0},h_{i};\pi_{i}),$
where $\mathcal{W}$ is the world model. Chaining $N$ primitives yields a full trajectory:
$\mathbf{x}_{1:T}=\text{Compose}(\{\mathcal{W}(\mathbf{x}_{t_{i-1}},h_{i};\pi_{i})\}_{i=1}^{N}),$
enabling long-horizon, high-level behaviors through explicit, interpretable composition. This design supports systematic generalization–novel sequences can be formed from known primitives, even if never seen during training.
@@ 3.2 Training Strategy
Sim-Real Hybrid Data Strategy In practice, we found that when training solely on real-world data, the model struggles to capture fine details of fast-moving parts such as the gripper, due to the reconstruction loss in the VAE. To enable the model to learn cleaner proprioceptive motion patterns, we introduced simulation data, including data from RLBench (James et al., 2020a) and LIBERO (Liu et al., 2024a), and adopt a sim-real hybrid data mixing strategy. For detailed data information, please refer to Appendix D.1.
This approach proves highly effective: the final model successfully combines the rich, complex textures from real-world data with the precise and clear kinematic motions from simulation, significantly enhancing the overall generation quality. Appendix I.3 Figure 12 shows a comprehensive comparison.
Three-Stage Finetuning Base Video Generation Model on Primitive Embodied Data We build upon the strong pretraining of large-scale video generation models. In this work, we introduce a three-stage fine-tuning strategy to adapt the model to primitive-based embodied tasks, balancing simulation-to-reality transfer and semantic alignment.
Stage 1: Simulation Pre-Finetuning. We first fine-tune the model on simulated data with a low number of epochs and early stopping, halting once generated videos exhibit plausible motion trends (e.g., arm movement toward an object). This rapid adaptation injects embodied semantics–such as robot dynamics, spatial interactions, and action semantics–into the model while avoiding premature convergence to simulation-specific modes.
Stage 2: Balanced Simulation-Real Mixing. We reduce the learning rate by half and train on a 1:1 mixture of real and simulated data. This stage encourages the model to align dynamics and appearance across domains, producing videos with clear, consistent actions and object structures in both settings.
Stage 3: Reality-Centric Refinement. Finally, we shift the data ratio to 80% real and 20% simulated, focusing the model on high-fidelity real-world generation while retaining the broad coverage of simulation. This stage emphasizes precise primitive-level temporal structure and fine-grained visual details.
For further training details (e.g., hyperparameters, augmentation, and evaluation criteria), please refer to Appendix F.2.
@@ 3.3 Causal Distillation and Acceleration: Real-Time On-the-Fly Future Frame Prediction
To address this, we adopt causal video generation via knowledge distillation, following CausVid (Yin et al., 2025). We train a student model to predict future frames in a strictly causal manner–each frame is generated based on past observations only, without access to future context. The student performs only 4 denoising steps per chunk (with 4 frames), enabling low-latency inference. To maintain generation quality under aggressive acceleration, we employ self-forcing–using the model’s own past predictions as input for future steps, closing the loop between prediction and conditioning.
@@ 4.1 Application 1: High-Quality Embodied Video Generation Facilitating Direct 6-DoF Trajectory Extraction
Prior methods often require a learning-based policy head or inverse kinematics (IK) module–after video generation, due to insufficient spatiotemporal precision. In contrast, we argue that our method is the first to make it practically feasible to directly extract end-effector 6-DoF trajectories from generated videos without any task-specific adaptation, as illustrated in Figure 4. This not only demonstrates the high spatiotemporal precision of our method, but also eliminates the need for additional learning or task-specific configurations.
@@
We apply Gen6D (Liu et al., 2022b), an off-the-shelf, zero-shot 6-DoF pose estimator, to the generated video frames to recover the full 6-DoF motion trajectory of the robot end-effector. This enables direct mapping from visual prediction to actionable control signals–bridging the gap between generative world models and real robot execution. Technical details of this extraction pipeline are provided in Appendix B. The experimental results are provided in Appendix B.2.

### v3

@@ Abstract
While video-generation-based embodied world models have gained increasing attention, their reliance on large-scale embodied interaction data remains a key bottleneck. The scarcity, difficulty of collection, and high dimensionality of embodied data fundamentally limit the alignment granularity between language and actions and exacerbate the challenge of long-horizon video generation–hindering generative models from achieving a "GPT moment" in the embodied domain. There is a naive observation: the diversity of embodied data far exceeds the relatively small space of possible primitive motions. Based on this insight, we propose Primitive Embodied World Models (PEWM), which restricts video generation to fixed shorter horizons, our approach 1) enables fine-grained alignment between linguistic concepts and visual representations of robotic actions, 2) reduces learning complexity, 3) improves data efficiency in embodied data collection, and 4) decreases inference latency. By equipping with a modular Vision-Language Model (VLM) planner and a Start-Goal heatmap Guidance mechanism (SGG), PEWM further enables flexible closed-loop control and supports compositional generalization of primitive-level policies over extended, complex tasks. Our framework leverages the spatiotemporal vision priors in video models and the semantic awareness of VLMs to bridge the gap between fine-grained physical interaction and high-level reasoning, paving the way toward scalable, interpretable, and general-purpose embodied intelligence.
@@ 1 Introduction
Figure 1: While densely distributed data can enable broad generalization, embodied data often suffer from sparsity (du2024compositional; xue2025demogen). The rightmost schematic highlights how organizing embodied data at the primitive level–along orthogonal dimensions such as action and object–supports compositional generalization even under limited data availability.
Embodied agents capable of planning and decision-making in complex environments rely heavily on internal representations of the world, commonly referred to as world models. Recent advances in pretrained video generation models–powered by self-supervised learning on vast internet-scale datasets–have demonstrated remarkable capabilities in synthesizing fine-grained, temporally coherent visual sequences from natural language descriptions (du2023learning).
These successes have inspired a growing body of work to leverage such models as world models for embodied agents, casting policy learning as a video generation problem (Ko2023Learning; du2023learning; yang2023learning; videoworldsimulators2024; black2024pi0visionlanguageactionflowmodel; xiang2024pandora; bruce2024genie; hu2024video; zhou2024robodreamerlearningcompositionalworld; dalal2024manipgen; liang2024dreamitaterealworldvisuomotorpolicy; reasoning2024arxiv; hu2024video; li2025unified; team2025aether; pertsch2025fast; zhen2025tesseract1; zhen2025tesseractlearning4dembodied; zhen2025tesseract; zhu2025uwm; hafner2025mastering): given a language-specified goal, the model generates a future video roll-out, from which actions are subsequently extracted. This paradigm offers compelling advantages in interpretability, generalization, and unified representation across diverse environments.
@@ 2.1 The Motivation and Definition of Primitive Embodied Data
“What can be said at all can be said clearly, and what we cannot talk about we must pass over in silence.”
—— Ludwig Josef Johann Wittgenstein
Collecting real-world embodied data is labor-intensive, and even large-scale datasets suffer from sparsity, high dimensionality, and limited generalization beyond minor scene variations (xue2025demogen; rt-1; open_x_embodiment_rt_x_2023; khazatsky2024droid; walke2023bridgedata; dalal2024manipgen), making fine-grained language-action alignment especially challenging. To overcome this, we propose organizing embodied experience at the level of primitives–short, language-grounded action units that serve as atomic building blocks for complex behaviors, enabling denser supervision and scalable compositional generalization.
Definition 2.1 (Primitive as Semantically Atomic Action Unit).
A primitive is a finite-duration embodied trajectory $p=(\mathbf{x}_{0:T},a_{0:T-1})$, paired with a natural language instruction $u$, such that:
  1. 1.
    Semantic atomicity: $u$ expresses a single, coherent manipulation intent that cannot be meaningfully decomposed into shorter language-grounded sub-instructions without loss of task-level meaning;
  2. 2.
    Temporal locality: $T$ is short enough (e.g., $\leq$ 2 seconds) to support high-fidelity video generation and precise 6-DoF trajectory extraction;
  3. 3.
    Generative feasibility: Given an initial observation and spatial guidance (e.g., start–goal heatmaps), a diffusion-based world model can reliably generate a plausible visual rollout of $p$.
Crucially, a primitive is NOT defined by mechanical simplicity, but by semantic indivisibility under language grounding. Thus, while “pick cup” qualifies, so does “arrange flowers”–provided it is executed as a unified intent within a short horizon and describable by a single instruction.
Building on this notion, we posit that complex tasks are inherently compositional:
Assumption 2.2 (Compositional Decomposability of Embodied Tasks).
Any long-horizon embodied task can be decomposed into a finite sequence of one or more primitives, each satisfying Definition 2.1.
This leads to a key structural insight:
Corollary 2.3 (Compact Primitive Template Basis).
The number of distinct primitive templates (e.g., “pick”, “open”, “arrange”) is vastly smaller than the total number of possible embodied trajectories, due to the combinatorial explosion of object, scene, and embodiment configurations.
This disparity implies that primitive-centric data organization is not merely convenient–it is fundamentally more efficient and scalable. Specifically, it enables:
  1. 1.
    High data efficiency: A single long-horizon demonstration yields multiple labeled primitives, dramatically increasing the effective sample count and mitigating data sparsity (Fig. 1);
  2. 2.
    Structured low-dimensional learning: Each primitive is short, visually coherent, and aligned with a single linguistic concept, facilitating fine-grained alignment between language and action in the video world model;
  3. 3.
    Direct action extraction: High-fidelity, short-horizon video generation enables zero-shot 6-DoF trajectory extraction (Appendix Sec. B), eliminating the need for task-specific policy heads;
  4. 4.
    Plug-and-play compositional generalization: A trained primitive world model becomes a reusable “skill library” that a high-level VLM planner can sequence on-the-fly (Fig. 2), supporting zero-shot execution of novel long-horizon tasks without retraining.
In sum, primitive embodied data provides a principled interface between symbolic reasoning (language) and continuous dynamics (video), turning the world model into a modular, interpretable, and scalable cerebellum for embodied intelligence.
2.2 Data Collection
@@
Enhancement in Efficiency, Density, and Quality  We construct a primitive-centric dataset $\mathcal{D}_{\text{prim}}=\{E_{i}\}$, where each episode $E_{i}$ is decomposed into primitives $\{\mathcal{P}^{k}\}_{k=1}^{N_{i}}$ with $\mathcal{P}^{k}=(\mathbf{x}_{t^{k}}^{\text{img}},\text{Instr}^{k},H^{k}_{s\to g},\mathbf{x}_{t^{k}+1:t^{k+1}-1}^{\text{img}})$. Here, $\mathbf{x}_{t^{k}}^{\text{img}}$ is the initial image, $\text{Instr}^{k}$ the instruction, $H^{k}_{s\to g}$ the start-goal heatmap, and future frames $\mathbf{x}_{t^{k}+1:t^{k+1}-1}^{\text{img}}$ are to be predicted.
@@
Full-Arm Shotting Facilitating Learning-Based Embodied World Simulator  Another notable design choice in $\mathcal{D}_{\text{prim}}$ is that the full robot arm–including the base and joint structures–is meant to be visible in the field of view, in contrast to cropped, end-effector-only frames common in prior datasets. This enables the diffusion model to internalize soft physical constraints such as reachability, joint limits, and spatial feasibility during prediction.
@@
The High Potential of Leveraging Action-Free Video Data for Scalable, Web-Scale Embodied Learning  It is worth noting that the LIBERO (liu2024libero) data we use not only includes the original dataset’s multi-view replays but also videos generated through OpenVLA rollouts as training data. This directly demonstrates our method’s ability to leverage any embodied video–whether from rollouts of other policies or from web-scale data–to improve model performance, highlighting one of the key advantages of using video as a universal representation.
@@ 3.1 Dual-Level Compositional Generalization
Implicit Compositionality of Diffusion Video Generation Diffusion models are capable of generating photo-realistic images that combine elements which likely do not appear together in the training set, demonstrating the ability to compositionally generalize (liu2022compositional; liang2024diffusion).
@@
We ground this compositional generalization in an Energy-Based Model (EBM) perspective. EBMs naturally support compositionality through additive energy decomposition: $E(\mathbf{x})=\sum_{i}E_{i}(\mathbf{x};\phi_{i}),$ where each term $E_{i}$ corresponds to a semantic factor (e.g., object, action, or spatial relation). Novel combinations arise by recombining known factors via energy minimization.
@@
Given the diverse and realistic co-occurrence of embodied concepts–such as shape, motion, and spatial relations–in our training data, the model learns rich semantic priors within each 32-frame primitive, enabling strong compositional generalization. Spatiotemporal coherence is inherited from large-scale video pretraining, a key advantage over static image-based methods. We treat each primitive as a distribution and use a heatmap as conditional input, formulating generation as: $\mathbf{x}_{1:T}^{\text{img}}\sim P(\mathbf{x}_{1:T}^{\text{img}}\mid H_{s\to g},\text{img}_{0})$. We adopt an I2V model as the world model $\mathcal{W}:(\mathbf{x}^{\text{img}}_{0},H_{s\to g})\mapsto\mathbf{x}^{\text{img}}_{1:T}$. The base model is DynamiCrafter (xing2024dynamicrafter), which is pre-trained on data without robotic arms. Our method performs well on robotic tasks, demonstrating effective zero-shot transfer. See Appendix F.1 for details.
Explicit Compositionality by Sequentially Combining Primitives By sequentially composing primitives, we treat the Primitive-Enabled World Model (PEWM) as a plug-and-play module for constructing complex behaviors. Each primitive $\pi_{i}\in\mathcal{P}$ represents a reusable, low-level dynamic policy that maps a current image $\mathbf{x}_{0}$ and a goal heatmap $h_{i}$ to a future video segment: $\mathbf{x}_{1:T_{i}}\sim\mathcal{W}(\mathbf{x}_{0},h_{i};\pi_{i})$, where $\mathcal{W}$ is the world model. Chaining $N$ primitives yields a full trajectory: $\mathbf{x}_{1:T}=\text{Compose}(\{\mathcal{W}(\mathbf{x}_{t_{i-1}},h_{i};\pi_{i})\}_{i=1}^{N})$, enabling long-horizon, high-level behaviors through explicit, interpretable composition. This design supports systematic generalization–novel sequences can be formed from known primitives, even if never seen during training.
@@ 3.2 Training Strategy
Sim-Real Hybrid Data Strategy In practice, we found that when training solely on real-world data, the model struggles to capture fine details of fast-moving parts such as the gripper, due to the reconstruction loss in the VAE. To enable the model to learn cleaner proprioceptive motion patterns, we introduced simulation data, including data from RLBench (james2020rlbench) and LIBERO (liu2024libero), and adopt a sim-real hybrid data mixing strategy. For detailed data information, please refer to Appendix D.1. This approach proves highly effective: the final model successfully combines the rich, complex textures from real-world data with the precise and clear kinematic motions from simulation, significantly enhancing the overall generation quality. Appendix I.3 Figure 12 shows a comprehensive comparison.
Three-Stage Finetuning Base Video Generation Model on Primitive Embodied Data We build upon the strong pretraining of large-scale video generation models. In this work, we introduce a three-stage fine-tuning strategy to adapt the model to primitive-based embodied tasks, balancing simulation-to-reality transfer and semantic alignment: Stage 1: Simulation Pre-Finetuning. We first fine-tune the model on simulated data with a low number of epochs and early stopping, halting once generated videos exhibit plausible motion trends (e.g., arm movement toward an object). This rapid adaptation injects embodied semantics–such as robot dynamics, spatial interactions, and action semantics–into the model while avoiding premature convergence to simulation-specific modes. Stage 2: Balanced Simulation-Real Mixing. We reduce the learning rate by half and train on a 1:1 mixture of real and simulated data. This stage encourages the model to align dynamics and appearance across domains, producing videos with clear, consistent actions and object structures in both settings. Stage 3: Reality-Centric Refinement. Finally, we shift the data ratio to 80% real and 20% simulated, focusing the model on high-fidelity real-world generation while retaining the broad coverage of simulation. This stage emphasizes precise primitive-level temporal structure and fine-grained visual details. For further training details (e.g., hyperparameters, augmentation, and evaluation criteria), please refer to Appendix F.2.
@@ 3.3 Causal Distillation and Acceleration: Real-Time On-the-Fly Future Frame Prediction
To address this, we adopt causal video generation via knowledge distillation, following Self Forcing (yin2025slow). We train a student model to predict future frames in a strictly causal manner–each frame is generated based on past observations only, without access to future context. The student performs only 4 denoising steps per chunk (with 4 frames), enabling low-latency inference. To maintain generation quality under aggressive acceleration, we employ self-forcing–using the model’s own past predictions as input for future steps, closing the loop between prediction and conditioning.
@@ 4.1 Application 1: High-Quality Embodied Video Generation Facilitating Direct 6-DoF Trajectory Extraction
Prior methods (du2023learning; zhen2025tesseract) often require a learning-based policy head or inverse kinematics (IK) module–after video generation, due to insufficient spatiotemporal precision. In contrast, we argue that our method is the first to make it practically feasible to directly extract end-effector 6-DoF trajectories from generated videos without any task-specific adaptation, as illustrated in Figure 4. This not only demonstrates the high spatiotemporal precision of our method, but also eliminates the need for additional learning or task-specific configurations.
@@
We apply Gen6D (gen6d), an off-the-shelf, zero-shot 6-DoF pose estimator, to the generated video frames to recover the full 6-DoF motion trajectory of the robot end-effector. This enables direct mapping from visual prediction to actionable control signals–bridging the gap between generative world models and real robot execution. Technical details of this extraction pipeline are provided in Appendix B. The experimental results are provided in Appendix B.2.